Rh=1/resistivity

what is the relation between resistivity and hall coefficient may I get the relation with proof


1
Expert's answer
2021-02-02T09:31:19-0500

Since hall coefficient is given by, "R_H =\frac{ V t}{IB}"

where V is the voltage, I is current, t is the thickness and B is the magnetic field.


By Ohm's Law, "R = \frac{V}{I} = \frac{\rho t}{A}" where "\rho" is the resistivity

Then, "R_H =\frac{ Rt}{B} = \frac{\rho t^2}{A B}"

where A is the area of cross section.